OpenStudy (anonymous):

can someone walk me through this? i need to solve for x, 1 over 4 (2x − 14) = 4

OpenStudy (yttrium):

Multiply the whole eq'n by 4(2x-14) . Then solve for combine similar terms

OpenStudy (anonymous):

8x - 56 =4 8x = 60 x = 7.5 !!!!

OpenStudy (anonymous):

Do you understand ?!

OpenStudy (anonymous):

4.2x-4.16=4 8x - 56 =4 8x=+56+4 8x = 60 x = 60/8 x = 7.5 :)It s soooooooooooooo easy !!!

OpenStudy (anonymous):

@Yttrium @E.ali yeah, thanks to the both of you!

OpenStudy (yttrium):

Hey!

OpenStudy (anonymous):

@:Yttium ,Amm . What is the over ?

OpenStudy (yttrium):

Wait. AHAHAHA

OpenStudy (yttrium):

@slumberousteenager is the question like this:\[(1/4) (2x-14) = 4\] if so, 1 over 4 (2x-14) = 4 Factor out 2, therefore x-7 = 8 x = 15

OpenStudy (anonymous):

Yes !!! I didnt know over ! Excuse me !

OpenStudy (anonymous):

But your answer is fall !